Spell checking software catches “non-word errors” that result in a string of letters that is not...

  1. Spell checking software catches “non-word errors” that result in a string of letters that is not a word, as when “the” is typed as “teh”. When undergraduates are asked to write a 250 – word essay (without spell checking) the number X of non-word errors has the following distribution:

Value of X

0

1

2

3

4

Probability

0.1

0.2

0.3

0.3

0.1

  1. Write the event “at least one non-word error in terms of X. What is the probability of this event?
  2. Describe the event X ≤ 2 in words. What is its probability? What is the probability

that X < 2?
Answer #1
X Probability
0 0.1
1 0.2
2 0.3
3 0.3
4 0.1

a) P(X ≥ 1) = P(1)+P(2)+P(3)+P(4)

= 0.2+0.3+0.3+0.1 = 0.9

b) X ≤ 2 means less than or equal to 2 non-word error .

P(X ≤ 2 ) = P(0)+P(1)+P(2)

= 0.1+0.2+0.3 = 0.6

P(X < 2) = P(0)+P(1) = 0.1+0.2 = 0.3